OpenStudy (anonymous):

Select all the pairs of ratios that form a proportion. A. 1/6, 4/20 B. 7/9, 28/36 C. 14/18, 21/27 D. 30/80, 6/18

OpenStudy (anonymous):

A. \[\frac{ 4 }{ 20 }=\frac{ 1 }{ 5 }\] B. \[\frac{ 28 }{36 }\div \frac{ 4 }{ 4 }=\frac{ 7 }{ 9 }\] C. \[\frac{ 14 }{18 }\div \frac{ 2 }{2 }=\frac{ 7 }{ 9 }\] \[\frac{ 21 }{ 27 }\div \frac{ 3 }{3 }=\frac{ 7 }{ 9 }\] D. \[\frac{ 30 }{ 80 }\div \frac{ 10 }{10 }=\frac{ 3 }{ 8 }\] \[\frac{ 6 }{18 }\div \frac{ 6 }{ 6 }=\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

So, B & C ?

OpenStudy (anonymous):

YES